# PINNfluence: Influence Functions for Physics-Informed Neural Networks

### v1

@@ A.1 Influence Functions
$\hat{\theta}_{\epsilon,\mathcal{X}^{+},\mathcal{X}^{-}}:=\operatorname*{arg\,min}_{\theta\in\Theta}\left[\frac{1}{N}\sum_{i=1}^{N}L(\boldsymbol{x}_{i};\theta)+\epsilon\sum_{i=1}^{m^{+}}L(\boldsymbol{x}_{i}^{+};\theta)-\epsilon\sum_{i=1}^{m^{-}}L(\boldsymbol{x}_{i}^{-};\theta)\right]\quad\text{,}$ (13)
where $\hat{\theta}_{0,\mathcal{X}^{+},\mathcal{X}^{-}}=\hat{\theta}$ is the optimized parameter for the training set $\mathcal{X}$. As we can multiply the inner function of Eq. 13 by any positive number without changing $\hat{\theta}_{\epsilon,\mathcal{X}^{+},\mathcal{X}^{-}}$, the optimal parameter for the training set $\mathcal{X}\cup\mathcal{X}^{+}\setminus\mathcal{X}^{-}$ is $\hat{\theta}_{1/N,\mathcal{X}^{+},\mathcal{X}^{-}}=\hat{\theta}_{\mathcal{X}^{+},\mathcal{X}^{-}}$. Therefore, we will consider
$\frac{f(\boldsymbol{z};\hat{\theta})-f(\boldsymbol{z};\hat{\theta}_{\epsilon,\mathcal{X}^{+},\mathcal{X}^{-}})}{\epsilon}$ $=-\left.\frac{\partial f(\boldsymbol{z};\hat{\theta}_{\epsilon,\mathcal{X}^{+},\mathcal{X}^{-}})}{\partial\epsilon}\right|_{\epsilon=0}+\mathcal{O}(\epsilon)$
$=-\nabla_{\hat{\theta}}f(\boldsymbol{z};\hat{\theta})\cdot\left.\frac{\partial\hat{\theta}_{\epsilon,\mathcal{X}^{+},\mathcal{X}^{-}}}{\partial\epsilon}\right|_{\epsilon=0}+\mathcal{O}(\epsilon)$
$=\nabla_{\hat{\theta}}f(\boldsymbol{z};\hat{\theta})\cdot H_{\hat{\theta}}^{-1}\cdot\nabla_{\hat{\theta}}\left(\sum_{i=1}^{m^{+}}L(\boldsymbol{x}_{i}^{+};\hat{\theta})-\sum_{i=1}^{m^{-}}L(\boldsymbol{x}_{i}^{-};\hat{\theta})\right)+\mathcal{O}(\epsilon)$ (14)
where we used Taylor’s theorem in the first step and Lemma 1 in the last step. ∎
Note, that the loss function will not be strictly convex in the most approaches, which implies, in particular, that it will not have a positive definite Hessian. However, in a progressing training, we can always restrict the possible parameter set $\Theta$ in a meaningful way, such that the loss function is strictly convex and having a positive definite Hessian.
Therefore, the influence function will approximate the change of a function $f(\boldsymbol{z};\hat{\theta})$, when retraining it on a different training set. Usually, one uses the loss itself $f(\boldsymbol{x}_{\text{test}};\hat{\theta})=L(\boldsymbol{x}_{\text{test}};\hat{\theta})$ on a test point for that function. However, we can also study the influence to physical observables, which are relevant for the considered problem. Since, we can also study the influence to mean values, we can therefore study also the influence to regions in our domain.
In analogous fashion, we can easily extend this framework to study the effect of upweighting or downweighting individual loss terms. This is in particular interesting for PINNs, since a wrong weighting of the loss terms, is known to reduce the accuracy of PINNs drastically [25].
Corollary 1.
Consider a total loss, which is built by several loss terms $\mathcal{L}_{j}$
@@
$\mathcal{L}(\mathcal{X};\theta)=\sum_{j=1}^{k}\mathcal{L}_{j}(\mathcal{X};\theta)=\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{k}l_{j}(\boldsymbol{x}_{i};\theta)\quad\text{.}$ (15)
The procedure of removing the loss terms $\mathcal{L}_{j}$ with $j\in K^{-}\subsetneq\{1,\ldots,k\}$ can be approximated by
$f(\boldsymbol{z};\hat{\theta})-f(\boldsymbol{z};\hat{\theta}_{K^{-}})=\frac{1}{N}\operatorname{Inf}_{f(\boldsymbol{z},\hat{\theta})}(K^{-})+\mathcal{O}(N^{-2})$ (16)
where $\hat{\theta}$ is the optimized parameter of the original model and $\hat{\theta}_{K^{-}}$ is the optimized parameter for the model with removed losses terms and
@@
$\operatorname{Inf}_{f(\boldsymbol{z},\hat{\theta})}(K^{-})\coloneq-\sum_{j\in K^{-}}\nabla_{\hat{\theta}}f(\boldsymbol{z};\hat{\theta})\cdot H_{\hat{\theta}}^{-1}\cdot\nabla_{\hat{\theta}}\mathcal{L}_{j}(\mathcal{X};\hat{\theta})\quad\text{.}$ (17)
Proof.
Analogous, to Proposition 1 we will define
$\hat{\theta}_{\epsilon,K^{-}}:=\operatorname*{arg\,min}_{\theta\in\Theta}\left[\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{k}l_{j}(\boldsymbol{x}_{i};\theta)-\epsilon\sum_{i=1}^{N}\sum_{j\in K^{-}}l_{j}(\boldsymbol{x}_{i};\theta)\right]$ (18)
as the optimal parameter, which will remove the loss terms $\mathcal{L}_{j}$ with $j\in K^{-}$ for $\epsilon=1/N$. The assertion follows with the analogue steps above. ∎
@@ A.3 Supplemental Figures
Figure 4: Trained models and target values with respective predictions: $u_{1}$, $u_{2}$, $p$, $|\!|\boldsymbol{u}|\!|$ (top to bottom).

### v2

@@ A.1 Influence Functions
$\operatorname{Inf}F[g_{0}+\epsilon\kappa]=-\left.\frac{dF(g_{0}+\epsilon\kappa)}{d\epsilon}\right|_{\epsilon=0}=\mathcal{H}_{\hat{\theta}}^{-1}\nabla_{\theta}\kappa(\hat{\theta})\quad\text{.}$ (11)
The corollary directly follows from the chain rule for differentiation. ∎
Note that the left-hand side of Eq. 9 is the Gâteaux derivative of $F$ at $g_{0}$ in the direction of $\kappa$, which provides a convenient mathematical setting to study influence functions for linear perturbations. We refer to [12] for the study of those relations.
In practice, we often use infinitesimal changes expressed by $\operatorname{Inf}f(F[g(\cdot,\epsilon)])$ to approximate some finite change in $f(F[g(\cdot,\epsilon)])$ for an $\epsilon>0$. From this perspective, the influence function is the linear approximation term of a Taylor series
$f(F[g(\cdot,\epsilon)])=f(F[g(\cdot,0)])-\epsilon\cdot\operatorname{Inf}f(F[g(\cdot,\cdot)])+r(\epsilon)\quad\text{,}$ (12)
where the remainder term $r(\epsilon)$ vanishes with
$\lim_{\epsilon\to 0}\frac{r(\epsilon)}{\epsilon^{2}}=0$ (13)
If, moreover, the functions $f$ and $g$ are analytic, then by Lemma 1, also $f(F[g(\cdot,\epsilon)])$ is analytic and the Taylor series converges uniformly for every compact subset of its convergence region, i.e. for any compact subset of $\mathbb{R}$ we have a constant $C>0$ such that the remainder is bounded there by
$\left|r(\epsilon)\right|\leq C\epsilon^{2}\quad\text{.}$ (14)
In the remaining part of this section, we will study concrete settings for influence functions, which appears in particular for PINNs, and give convenient reformulations of the statement of Theorem 1.
Adding & removing training points:
Let $\mathcal{X}$ be a finite, non-empty set of training points and consider a finite set of points $\mathcal{X}^{+}$, which we would like to add to the training set, as well as a set of points $\mathcal{X}^{-}\subsetneq\mathcal{X}$, which we would like to remove from the training set. Moreover, let $L(\boldsymbol{x},\cdot):\Theta\to\mathbb{R}$ be a twice continuously differentiable loss function assigning a value to each (potential) training point of $\mathcal{X}\cup\mathcal{X}^{+}$. The empirical risk for our original dataset is given by $g_{0}(\theta)=\frac{1}{N}\sum_{\boldsymbol{x}\in\mathcal{X}}L(\boldsymbol{x};\theta)$ and let $\kappa(\theta)=\sum_{\boldsymbol{x}\in\mathcal{X}^{+}}L(\boldsymbol{x};\theta)-\sum_{\boldsymbol{x}\in\mathcal{X}^{-}}L(\boldsymbol{x};\theta)$ be the perturbation in risk corresponding to the change in the training set. $N$ denotes the size of training points in $\mathcal{X}$. Assume that there exists a stationary point [^3] Thus, in practice, we assume that our training has progressed to the stage where it has reached a stationary point in the loss landscape. (a global minimum, respectively) $\hat{\theta}=F[g_{0}]$ inside $\Theta$ and let $F$ be the functional which maps to those stationary points. In addition, let $f(\boldsymbol{z},\cdot):\Theta\to\mathbb{R}^{m}$ be an arbitrary continuously differentiable function (e.g. the loss or the prediction), possibly with a further dependency expressed by a variable $\boldsymbol{z}$ (e.g. a test point).
According to Theorem 1, we can then approximate counterfactual effect of training on a perturbed training set $\mathcal{X}^{\prime}=(\mathcal{X}\cup\mathcal{X}^{+})\setminus\mathcal{X}^{-}$ by
@@
$f(\boldsymbol{z};\hat{\theta})-f(\boldsymbol{z};\hat{\theta}^{\prime})=\frac{1}{N}\operatorname{Inf}_{f(\boldsymbol{z};\hat{\theta})}(\mathcal{X}^{+},\mathcal{X}^{-})+r\!\left(N^{-2}\right)\quad\text{,}$ (15)
where we introduce the shorthand notation
$\operatorname{Inf}_{f(\boldsymbol{z};\hat{\theta})}(\mathcal{X}^{+},\mathcal{X}^{-})$ :-$\coloneq\operatorname{Inf}f(\boldsymbol{z};F[g_{0}+\epsilon\kappa]):$
@@
$=\nabla_{\theta}f(\boldsymbol{z};\hat{\theta})\cdot\mathcal{H}_{\hat{\theta}}^{-1}\cdot\nabla_{\theta}\left(\sum_{\boldsymbol{x}\in\mathcal{X}^{+}}L(\boldsymbol{x};\hat{\theta})-\sum_{\boldsymbol{x}\in\mathcal{X}^{-}}L(\boldsymbol{x};\hat{\theta})\right)\quad\text{,}$ (16)
which is simply the signed sum of the influences of the individual training points due to the linearity of Influence Functions.
Removing a complete loss term:
Moreover, we can also remove a complete loss term and study the influence of this operation to the training. For this, consider a total loss as a sum of $k$ loss terms
$g_{0}(\theta)=\mathcal{L}(\mathcal{X};\theta)=\sum_{j=1}^{k}\mathcal{L}_{j}(\mathcal{X};\theta)=\frac{1}{N}\sum_{\boldsymbol{x}\in\mathcal{X}}\sum_{j=1}^{k}L_{j}(\boldsymbol{x};\theta)$ (17)
[…]
with
$\operatorname{Inf}f(\boldsymbol{z};F[g_{0}+N^{-1}\kappa])=-\nabla_{\theta}f(\boldsymbol{z};\hat{\theta})\cdot\mathcal{H}_{\hat{\theta}}^{-1}\cdot\nabla_{\theta}\sum_{\boldsymbol{x}\in\mathcal{X}}L_{i}(\boldsymbol{x};\hat{\theta})\quad\text{.}$ (19)
Influence of loss weights:
In analogous fashion, we can easily extend this framework to study the effect of upweighting or downweighting individual loss terms. This is in particular interesting for PINNs because a wrong weighting of the loss terms is known to reduce the accuracy of PINNs drastically [31].
Thus, assume that for some weights $w_{j}>0$ we consider a total loss in the form of
$g_{0}(\theta)=\mathcal{L}(\mathcal{X};\theta)=\sum_{j=1}^{k}w_{j}\mathcal{L}_{j}(\mathcal{X};\theta)$ (20)
where we now perturb the weight $w_{i}$ with $\kappa(\theta)=w_{i}\mathcal{L}_{i}(\mathcal{X};\theta)$. The influence of changing the weight $w_{i}$ is then given by
$\operatorname{Inf}f(\boldsymbol{z};F[g_{0}+\epsilon\kappa])=w_{i}\nabla_{\theta}f(\boldsymbol{z};\hat{\theta})\cdot\mathcal{H}_{\hat{\theta}}^{-1}\cdot\nabla_{\theta}\mathcal{L}_{i}(\mathcal{X},\hat{\theta})\quad\text{.}$ (21)
In comparison to the two previous discussed cases, we can also consider infinitesimal changes of the weight so that we can omit a discussion of the error bounds.
Influence of parameters in the PDE:
As a last option to use influence functions for PINNs, we want to discuss the impact of parameters in the PDE. Therefore, we will consider a basic objective of the form
$g(\theta,\rho)=\frac{1}{N}\sum_{\boldsymbol{x}\in\mathcal{X}}L(\boldsymbol{x};\rho;\theta)$ (22)
with some additional parameter $\rho\in\mathbb{R}$, and we will assume that $\mathcal{L}$ is also continuously differentiable w.r.t. $\rho$. To apply Lemma 1, we will assume, that there is an optimal or stationary parameter $\hat{\theta}=F[g(\cdot,\rho_{0})]$ at a certain value of the parameter $\rho_{0}$ we are interested in. Then, the influence of the parameter $\rho$ to the optimal parameters $\hat{\theta}$ at a certain point $\rho_{0}$ is given by
$\operatorname{Inf}F[g]$ $=\lim_{\epsilon\to 0}\frac{F[g(\theta,\rho_{0})]-F[g(\theta,\rho_{0}+\epsilon)]}{\epsilon}=-\left.\frac{dF[g(\theta,\rho)]}{d\rho}\right|_{\rho=\rho_{0}}$
$=\mathcal{H}_{\theta}^{-1}(\hat{\theta},\rho_{0})\cdot\nabla_{\theta}\frac{\partial}{\partial\rho}g(\hat{\theta},\rho_{0})\quad\text{.}$ (23)
@@ A.3 Supplemental Figures
Figure 4: Trained models and target values with respective predictions for $u_{1}$, $u_{2}$, $p$, $|\!|\boldsymbol{u}|\!|$ (top to bottom).